solve the undetermined coefficient y''+7y' =42x2+5x+1

1
Expert's answer
2021-06-01T06:25:44-0400

Let us solve the differential equation "y''+7y' =42x^2+5x+1."


The characteristic equation "k^2+7k=0" has the solutions "k_1=0" and "k_2=-7." Therefore, the general solution of the differential equation is of the form "y(x)=C_1+C_2e^{-7x}+y_p." Since "k=0"

is a simple root of the characteristic equation, "y_p=x(ax^2+bx+c)=ax^3+bx^2+cx."


"y_p'=3ax^2+2bx+c,\ \ y_p''=6ax+2b."


It follows that "6ax+2b+7(3ax^2+2bx+c)=21ax^2+(6a+14b)x+(2b+7c)=42x^2+5x+1."

Consequently, "21a=42, \ 6a+14b=5, \ 2b+7c=1." Therefore, "a=2,\ b=-\frac{1}{2},\ c=\frac{2}{7}."


We concude that the general solution of the differential equation is of the form "y(x)=C_1+C_2e^{-7x}+2x^3-\frac{1}{2}x^2+\frac{2}{7}x."